# Can Knowledge Graphs Make Large Language Models More Trustworthy? An Empirical Study over Open-ended Question Answering

### v2

@@ A.2 Query Construction
Specifically, we consider five angles to measure the quality of the generated queries: (1) Naturalness: assessing how fluid and human-like the query sounds; (2) Relevance: measuring whether the query pertains directly to the entity and the context provided; (3) Specificity: determining the level of detail and granularity included in the query, ensuring it is not too broad or vague; (4) Novelty: evaluating the uniqueness of the query, ensuring it is not just a repetitive or common question; (5) Actionability: gauging whether the query prompts clear, definite answers or actions that are feasible within the given context. Each of these angles contributes to a holistic evaluation of the query’s effectiveness and relevance in real-world applications.
We have three evaluators participating in the manual assessment of query quality. All of the evaluators are computer science majors with fluent English skills. As the evaluation centers on various linguistic metrics such as naturalness, relevance, specificity, novelty, and actionability, we only require the evaluators to possess a fundamental understanding of English without restricting their majors. We calculate the Pearson correlation coefficients between human and LLM scores as shown in Table 3. It shows that as the rounds progress, agreement between humans and LLMs increases, suggesting that iterative feedback improves alignment between human annotation and LLM responses.
@@
Table 4: Cohen’s Kappa coefficient for various metrics.
@@
A.3 Personalized PageRank (PPR)
@@
A.4 Prize-Cost-based Path Retrieval
The path-retrieval method is designed to construct and evaluate paths in a graph based on predefined prize assignments and cost allocations. The objective is to form sequences of nodes and edges, represented as $\mathcal{P}=\{v_{1},e_{1},v_{2},\dots,e_{n-1},v_{n}\}$, that maximize the overall score and minimize the costs. To efficiently manage the exploration of potential paths, we utilize a priority queue, a data structure that allows paths to be organized based on their scores, ensuring that the highest-scoring paths are processed first. The method starts by picking a number of starting nodes with high prizes. We then expand these starting points by exploring neighboring nodes. For each neighbor, the method calculates a new score. This score is the sum of the neighbor’s prize and the edge’s prize minus the edge’s cost. If this neighbor hasn’t been visited before, which helps avoid looping, the algorithm adds this neighbor to the path. This new path is then added to the priority queue. This expansion keeps going until the path reaches a maximum length or can’t be extended further. The algorithm keeps track of paths already explored to avoid repetition and ensure paths don’t loop back on themselves. When no more paths can be added or the priority queue is empty, the algorithm sorts the paths by their scores from highest to lowest.
@@ A.5 KG Similarity Metrics
To measure how much the perturbed KGs are deviated from the original KG used in §2. We re-present the according metrics from (Raman et al., 2020) below:
@@
Aggregated Triple Score (ATS)  ATS measures semantic similarity between two KGs. Let $s_{\mathcal{G}}$ be an edge (triple) scoring function, such that $s_{\mathcal{G}}(e_{1},r,e_{2})$ measures how likely edge $(e_{1},r,e_{2})$ is to exist in $\mathcal{G}$. Also, assume $s_{\mathcal{G}}$ has been pre-trained on $\mathcal{G}$ for link prediction. Then, ATS is defined as $f_{\text{ATS}}(\mathcal{G},\mathcal{G}^{{}^{\prime}})=\frac{1}{|\mathcal{T}^{{}^{\prime}}|}\sum_{(e_{1},r,e_{2})\in\mathcal{T}^{{}^{\prime}}}s_{\mathcal{G}}(e_{1},r,e_{2})\in[0,1]$, which denotes the mean $s_{\mathcal{G}}$ score across all edges in $\mathcal{G^{\prime}}$. Intuitively, if a high percentage of edges in $\mathcal{G^{\prime}}$ are also likely to exist in $\mathcal{G}$ (i.e., high ATS), then we say that $\mathcal{G^{\prime}}$ and $\mathcal{G}$ have high semantic similarity. $s_{\mathcal{G}}$ is task-specific, as KGs from different tasks may differ greatly in semantics. We use the $s_{\mathcal{G}}$ from (Li et al., 2016); while ATS captures semantic KG differences, it is not sensitive to KG connectivity structure. Note that $f_{\text{ATS}}(\mathcal{G},\mathcal{G})$ may not equal 1, since $s_{\mathcal{G}}$ may not perfectly generalize to KGs beyond those it was trained on.
@@
Similarity in Clustering Coefficient Distribution (SC2D)  SC2D measures structural similarity between two KGs and is derived from the local clustering coefficient (Saramäki et al., 2007; Onnela et al., 2005; Fagiolo, 2007). For a given entity in $\mathcal{G}$ (treated here as undirected), the local clustering coefficient is the fraction of possible triangles through the entity that exist (i.e., how tightly the entity’s neighbors cluster around it). For entity $e_{i}\in\mathcal{E}$, the local clustering coefficient is defined as $c_{i}=2\text{Tri}(e_{i})/(\text{deg}(e_{i})(\text{deg}(e_{i})-1))$, where $\text{Tri}(e_{i})$ is the number of triangles through $e_{i}$, and $\text{deg}(e_{i})$ is the degree of $e_{i}$. For each relation $r\in\mathcal{R}$, let $\mathcal{G}^{r}$ be the subgraph of $\mathcal{G}$ consisting of all edges in $\mathcal{T}$ with $r$ That is, $\mathcal{G}^{r}=(\mathcal{E},r,\mathcal{T}^{{}^{\prime}})$, where $\mathcal{T}^{{}^{\prime}}=\{(e,r,e^{\prime})\hskip 2.84526pt|\hskip 2.84526pte,e^{\prime}\in\mathcal{E}\}$. Let $\mathbf{c}^{r}$ denote the $|\mathcal{E}|$-dimensional clustering coefficient vector for $\mathcal{G}^{r}$, where the $i$th element of $\mathbf{c}^{r}$ is $c_{i}$. Then, the mean clustering coefficient vectors for $\mathcal{G}$ and $\mathcal{G}^{\prime}$ are $\mathbf{c}_{o}=\frac{1}{|\mathcal{R}|}\sum_{r\in\mathcal{R}}\mathbf{c}^{r}$ and $\mathbf{c}_{p}=\frac{1}{|\mathcal{R}^{\prime}|}\sum_{r\in\mathcal{R}^{\prime}}\mathbf{c}^{r}$, respectively. SC2D is defined as $f_{\text{SC2D}}(\mathcal{G},\mathcal{G}^{{}^{\prime}})=1-\frac{\lVert\mathbf{c}_{o}-\mathbf{c}_{p}\rVert_{2}}{\lVert\mathbf{c}_{o}-\mathbf{c}_{p}\rVert_{2}+1}\in[0,1]$, with higher value indicating higher similarity.
@@
Similarity in Degree Distribution (SD2)  SD2 also measures structural similarity between two KGs, while addressing SC2D’s ineffectiveness when the KGs’ entities have tiny local clustering coefficients (e.g., the item KG used by recommender systems is roughly bipartite). In such cases, SC2D is always close to one regardless of the perturbation method, thus rendering SC2D useless. Let $\mathbf{d}^{r}$ denote the $|\mathcal{E}|$-dimensional degree vector for $\mathcal{G}^{r}$, where the $i$th element of $\mathbf{d}^{r}$ is $\text{deg}(e_{i})$. Then, the mean degree vectors for $\mathcal{G}$ and $\mathcal{G}^{\prime}$ are $\mathbf{d}_{o}=\frac{1}{|\mathcal{R}|}\sum_{r\in\mathcal{R}}\mathbf{d}^{r}$ and $\mathbf{d}_{p}=\frac{1}{|\mathcal{R}^{\prime}|}\sum_{r\in\mathcal{R}^{\prime}}\mathbf{d}^{r}$, respectively. SD2 is defined as $f_{\text{SD2}}(\mathcal{G},\mathcal{G}^{{}^{\prime}})=1-\frac{\lVert\mathbf{d}_{o}-\mathbf{d}_{p}\rVert_{2}}{\lVert\mathbf{d}_{o}-\mathbf{d}_{p}\rVert_{2}+1}\in[0,1]$, with higher value indicating higher similarity.
Appendix B Related Work
@@
Researchers employ diverse strategies to augment the LLMs by integrating external KGs (Sui et al., 2024; He et al., 2024b). For example, KAPING (Baek et al., 2023b) matches entities in questions to retrieve related triples from knowledge graphs for zero-shot question answering. Wu et al. (2023) finds that converting these triples into textualized statements can further enhance LLM performance. StructGPT (Jiang et al., 2023b) propose to convert user query into structured formats (e.g., SPARQL) for information extraction from KGs. Following the succuess of internal reasoning-enhancement methods like Chain-of-thoughts (CoT) (Wei et al., 2022), Reflexion (Shinn et al., 2024), and Tree-of-thoughts (ToT), He et al. (2022) propose “rethinking with retrieval” to use decomposed reasoning steps from CoT prompting to retrieve external knowledge, leading to more accurate and faithful explanations. IR-CoT (Trivedi et al., 2022) interleaves the generation of CoT with knowledge retrieval from corresponding KGs, iteratively guiding both retrieval and reasoning for multi-step questions. MindMap (Wen et al., 2023) introduce a plug-and-play approach to evoke graph-of-thoughts reasoning in LLMs. Similarly, RoG (Luo et al., 2023b) use KGs to create faithful reasoning paths based on various relations, enabling interpretable reasoning in LLMs.
@@
Appendix C Prompt List
@@
C.1 Knowledge-Augmented Generation

### v3

@@ A.1.1 Human-in-the-loop for instruction optimization
A.1.2 Metrics for generated queries
We consider five metrics to measure the quality of the generated queries: (1) Naturalness: assessing how fluid and human-like the query sounds; (2) Relevance: measuring whether the query pertains directly to the entity and the context provided; (3) Specificity: determining the level of detail and granularity included in the query, ensuring it is not too broad or vague; (4) Novelty: evaluating the uniqueness of the query, ensuring it is not just a repetitive or common question; (5) Actionability: gauging whether the query prompts clear, definite answers or actions that are feasible within the given context. Each of these angles contributes to a holistic evaluation of the query’s effectiveness and relevance in real-world applications.
A.1.3 Verifying human-in-the-loop
For the human-label scores $s_{\mathbf{human}}$ collection, we have three evaluators participating in the manual assessment of query quality. All of the evaluators are computer science majors with fluent English skills. As the evaluation centers on various linguistic metrics such as naturalness, relevance, specificity, novelty, and actionability, we only require the evaluators to possess a fundamental understanding of English without restricting their majors. We calculate the Pearson correlation coefficients between human and LLM scores as shown in Table 3. It shows that as the rounds progress, agreement between humans and LLMs increases, suggesting that iterative feedback improves alignment between human annotation and LLM responses.
@@ A.1.3 Verifying human-in-the-loop
Table 4: Cohen’s Kappa coefficient for various metrics.
A.2 Personalized PageRank (PPR)
@@
Setting Context Relevance Comprehensiveness Correctness Empowerment SAFE FActScore
OKGQA (subgraphs) $75.70\%\pm 0.44\%$ $71.51\%\pm 0.83\%$ $66.43\%\pm 0.76\%$ $69.60\%\pm 0.65\%$ $88.71\%\pm 0.72\%$ $70.12\%\pm 0.87\%$
+ Multi-lingual context $75.14\%\pm 0.33\%$ $72.32\%\pm 0.19\%$ $66.72\%\pm 0.74\%$ $70.32\%\pm 0.57\%$ $90.32\%\pm 0.48\%$ $72.83\%\pm 0.93\%$
Table 5: Comparison of GPT-4o-mini Performance Using Monolingual and Multilingual Subgraphs
A.3 Prize-Cost-based Path Retrieval
In this section, we detail the path-retrieval method used in §3.2. It is designed to construct and evaluate paths in a graph based on predefined prize assignments and cost allocations. The objective is to form sequences of nodes and edges, represented as $\mathcal{P}=\{v_{1},e_{1},v_{2},\dots,e_{n-1},v_{n}\}$, that maximize the overall score and minimize the costs. To efficiently manage the exploration of potential paths, we utilize a priority queue, a data structure that allows paths to be organized based on their scores, ensuring that the highest-scoring paths are processed first. The method starts by picking a number of starting nodes with high prizes. We then expand these starting points by exploring neighboring nodes. For each neighbor, the method calculates a new score. This score is the sum of the neighbor’s prize and the edge’s prize minus the edge’s cost. If this neighbor hasn’t been visited before, which helps avoid looping, the algorithm adds this neighbor to the path. This new path is then added to the priority queue. This expansion keeps going until the path reaches a maximum length or can’t be extended further. The algorithm keeps track of paths already explored to avoid repetition and ensure paths don’t loop back on themselves. When no more paths can be added or the priority queue is empty, the algorithm sorts the paths by their scores from highest to lowest.
A.4 LLM Evaluation Clarity
To address the concern regarding potential self-enhancement bias in LLM-as-evaluator frameworks, we have conducted extensive validation of our evaluation approach. In specific, we randomly sample 100 questions and evaluated them using three different LLMs (gpt-4o-mini, llama-3.1-8b-instruct, and gemma-2-9b-it). We measured inter-model agreement using Cohen’s Kappa as shown in Table 6, which showed substantial to almost perfect consistency. This indicates that the evaluation results are consistent across different LLMs, even when the model generating the responses is not the same as the one evaluating them (e.g., using gpt-4o-mini for generation and llama-3.1-8b-instruct for evaluation). These findings confirm that the evaluation is robust and independent of the specific LLM used as the evaluator.
Metric LLM 1 & 2 LLM 1 & 3 LLM 2 & 3
G-Eval 0.84 0.81 0.82
FactScore 0.78 0.74 0.78
SAFE 0.74 0.70 0.72
Table 6: Cohen’s Kappa coefficient for different LLM pair comparisons. For the G-Eval, we use the average score of four sub-metrics for better readability. LLM 1: gpt-4o-mini; LLM 2: llama-3.1-8b-instruct; LLM 3: gemma-2-9b-it)
In addition, we also collect human evaluations for these 100 samples. Three experts annotators rate each anonymized response on context relevance, comprehensiveness, correctness and empowerment using a 1-5 Likert scale. The average human ratings are computed and compared with automated scores using G-Eval. The Pearson’s correlation yields a score of 0.78, indicating strong alignment between human judgment and LLM-based evaluation. Combined with the inter-model agreement shown in Table 6, these results confirm that our evaluation is robust, consistent, and largely independent of the specific LLM used as the evaluator.
@@ A.5 KG Similarity Metrics
In this section, we introduce the metrics used in §2 to measure the deviation of the perturbed KGs from the original KG. These metrics are adapted from  Raman et al. (2020) as presented below. ATS is mainly used to measure the semantic similarity between two KGs, while SC2D and SD2 are used to measure the structural similarity.
Aggregated Triple Score (ATS):
@@
ATS measures semantic similarity between two KGs. Let $s_{\mathcal{G}}$ be an edge (triple) scoring function, such that $s_{\mathcal{G}}(e_{1},r,e_{2})$ measures how likely edge $(e_{1},r,e_{2})$ is to exist in $\mathcal{G}$. Also, assume $s_{\mathcal{G}}$ has been pre-trained on $\mathcal{G}$ for link prediction. Then, ATS is defined as $f_{\text{ATS}}(\mathcal{G},\mathcal{G}^{{}^{\prime}})=\frac{1}{|\mathcal{T}^{{}^{\prime}}|}\sum_{(e_{1},r,e_{2})\in\mathcal{T}^{{}^{\prime}}}s_{\mathcal{G}}(e_{1},r,e_{2})\in[0,1]$, which denotes the mean $s_{\mathcal{G}}$ score across all edges in $\mathcal{G^{\prime}}$. Intuitively, if a high percentage of edges in $\mathcal{G^{\prime}}$ are also likely to exist in $\mathcal{G}$ (i.e., high ATS), then we say that $\mathcal{G^{\prime}}$ and $\mathcal{G}$ have high semantic similarity. $s_{\mathcal{G}}$ is task-specific, as KGs from different tasks may differ greatly in semantics. We use the $s_{\mathcal{G}}$ from (Li et al., 2016); while ATS captures semantic KG differences, it is not sensitive to KG connectivity structure. Note that $f_{\text{ATS}}(\mathcal{G},\mathcal{G})$ may not equal 1, since $s_{\mathcal{G}}$ may not perfectly generalize to KGs beyond those it was trained on.
Similarity in Clustering Coefficient Distribution (SC2D):
@@
SC2D measures structural similarity between two KGs and is derived from the local clustering coefficient (Saramäki et al., 2007; Onnela et al., 2005; Fagiolo, 2007). For a given entity in $\mathcal{G}$ (treated here as undirected), the local clustering coefficient is the fraction of possible triangles through the entity that exist (i.e., how tightly the entity’s neighbors cluster around it). For entity $e_{i}\in\mathcal{E}$, the local clustering coefficient is defined as $c_{i}=2\text{Tri}(e_{i})/(\text{deg}(e_{i})(\text{deg}(e_{i})-1))$, where $\text{Tri}(e_{i})$ is the number of triangles through $e_{i}$, and $\text{deg}(e_{i})$ is the degree of $e_{i}$. For each relation $r\in\mathcal{R}$, let $\mathcal{G}^{r}$ be the subgraph of $\mathcal{G}$ consisting of all edges in $\mathcal{T}$ with $r$ That is, $\mathcal{G}^{r}=(\mathcal{E},r,\mathcal{T}^{{}^{\prime}})$, where $\mathcal{T}^{{}^{\prime}}=\{(e,r,e^{\prime})\hskip 2.84526pt|\hskip 2.84526pte,e^{\prime}\in\mathcal{E}\}$. Let $\mathbf{c}^{r}$ denote the $|\mathcal{E}|$-dimensional clustering coefficient vector for $\mathcal{G}^{r}$, where the $i$th element of $\mathbf{c}^{r}$ is $c_{i}$. Then, the mean clustering coefficient vectors for $\mathcal{G}$ and $\mathcal{G}^{\prime}$ are $\mathbf{c}_{o}=\frac{1}{|\mathcal{R}|}\sum_{r\in\mathcal{R}}\mathbf{c}^{r}$ and $\mathbf{c}_{p}=\frac{1}{|\mathcal{R}^{\prime}|}\sum_{r\in\mathcal{R}^{\prime}}\mathbf{c}^{r}$, respectively. SC2D is defined as $f_{\text{SC2D}}(\mathcal{G},\mathcal{G}^{{}^{\prime}})=1-\frac{\lVert\mathbf{c}_{o}-\mathbf{c}_{p}\rVert_{2}}{\lVert\mathbf{c}_{o}-\mathbf{c}_{p}\rVert_{2}+1}\in[0,1]$, with higher value indicating higher similarity.
Similarity in Degree Distribution (SD2):
@@
SD2 also measures structural similarity between two KGs, while addressing SC2D’s ineffectiveness when the KGs’ entities have tiny local clustering coefficients (e.g., the item KG used by recommender systems is roughly bipartite). In such cases, SC2D is always close to one regardless of the perturbation method, thus rendering SC2D useless. Let $\mathbf{d}^{r}$ denote the $|\mathcal{E}|$-dimensional degree vector for $\mathcal{G}^{r}$, where the $i$th element of $\mathbf{d}^{r}$ is $\text{deg}(e_{i})$. Then, the mean degree vectors for $\mathcal{G}$ and $\mathcal{G}^{\prime}$ are $\mathbf{d}_{o}=\frac{1}{|\mathcal{R}|}\sum_{r\in\mathcal{R}}\mathbf{d}^{r}$ and $\mathbf{d}_{p}=\frac{1}{|\mathcal{R}^{\prime}|}\sum_{r\in\mathcal{R}^{\prime}}\mathbf{d}^{r}$, respectively. SD2 is defined as $f_{\text{SD2}}(\mathcal{G},\mathcal{G}^{{}^{\prime}})=1-\frac{\lVert\mathbf{d}_{o}-\mathbf{d}_{p}\rVert_{2}}{\lVert\mathbf{d}_{o}-\mathbf{d}_{p}\rVert_{2}+1}\in[0,1]$, with higher value indicating higher similarity.
Appendix B Extension of OKGQA
In this section, we extend our benchmark by incorporating multilingual context and validating our query generation against DBpedia’s structure. We first introduce the multilingual setup of our dataset anc compare the performance of multilingual subgraphs with the monolingual subgraphs (§B.1). We then analyze the relationship between generated queries and DBpedia by examining query generation, entity/relation coverage, and subgraph alignment (§B.2). We also compare OKGQA with the existing widely used benchmarks in Table 7.
B.1 Multilingual Setup of OKGQA
KGs typically include entities and relations in multiple languages, providing a richer context that can benefit our OKGQA setting. In this experiment, we investigate whether incorporating multilingual context improves overall performance. Specifically, we randomly sample 300 queries from our dataset and generate subgraphs that include multilingual entities and relations from DBpedia. We then apply PPR consistent with our original method in §2.1 to reduce the KG size. For this multilingual setting, we consider five languages—Greek, Polish, Portuguese, Spanish, and English—which cover the majority of entities in DBpedia. We compare the performance of GPT-4o-mini using the new multilingual subgraphs against the original monolingual subgraphs, as shown in Table 5. Our findings indicate that including multilingual context generally leads to better performance across multiple metrics. Intuitively, this additional multi-lingual context may provides more knowledge from different perspectives (which could provide more context, but also may requires more techniques for handle challenges like duplicates across languages) and also provide another way to validate the factuality of the resources stored in the KGs (which can provide more authenticity through cross validation from different languages).
B.2 Generated Query-DBpedia Alignment
We analyze the alignment between our generated queries and DBpedia along three dimensions: query generation, entity/relation coverage, and subgraph alignment as follows:
Query Generation:
Each query is directly generated from DBpedia entities and their relationships. For example, when asking about Microsoft’s founder, we first confirm that both “Microsoft” and “Bill Gates” exist in DBpedia and are connected by the founded_by relation, ensuring that our queries are firmly grounded in the knowledge graph.
Entity and Relation Coverage:
Our analysis indicates that:
  • 92% entities mentioned in the queries can be detected from DBpedia entities.
  • 87% queries have complete relation paths connecting the relevant entities from DBPedia.
  • Entities/relations mentioned in queries cover 72% of DBpedia’s most common entities/predicates and span diverse entity types (e.g., Person, Organization, and Event).
Subgraph Alignment:
We evaluate the structure of the sampled subgraphs for each query and find that:
  • 75% of the queries retrieve subgraphs within 3–4 hops, which aligns with the typical depth for DBpedia reasoning tasks.
  • On average, each subgraph contains 48 nodes and 152 edges, with an average node degree of 3.17 and a clustering coefficient of 0.69, which also aligns with the property of DBPedia.
These statistics support that our dataset accurately reflects DBpedia’s structure, ensuring both authenticity and complexity in the generated queries.
Dataset # Questions Question Type Focus Areas Source of Questions Knowledge Base Hallucination Detection Unreliable KG
OKGQA 850 / 2,050 Open-ended Evaluating hallucination and reasoning capabilities in LLMs when augmented with Knowledge Graphs; diverse queries requiring complex reasoning Curated DBPedia
[…]
WebQuestions 5,810 Factoid Questions derived from Google Suggest queries, focusing on simple factual information User queries Freebase
[…]
ComplexWebQuestions 34,689 Multi-hop Factoid Extends WebQuestions with more complex, multi-hop questions requiring compositional reasoning User queries Freebase
[…]
GrailQA 64,331 Varied Factoid Evaluates generalization in KBQA with questions requiring different levels of reasoning Crowdsourced Freebase
Table 7: Comparison of OKGQA with existing benchmarks along with their question types, focus areas, and additional properties.
Appendix C Related Work
@@
Researchers employ diverse strategies to augment the LLMs by integrating external KGs (Sui et al., 2024; He et al., 2024b). For example, KAPING (Baek et al., 2023b) matches entities in questions to retrieve related triples from knowledge graphs for zero-shot question answering. Wu et al. (2023) finds that converting these triples into textualized statements can further enhance LLM performance. StructGPT (Jiang et al., 2023b) propose to convert user query into structured formats (e.g., SPARQL) for information extraction from KGs. Following the succuess of internal reasoning-enhancement methods like Chain-of-thoughts (CoT) (Wei et al., 2022), Reflexion (Shinn et al., 2024), and Tree-of-thoughts (ToT), He et al. (2022) propose “rethinking with retrieval” to use decomposed reasoning steps from CoT prompting to retrieve external knowledge, leading to more accurate and faithful explanations. IR-CoT (Trivedi et al., 2022b) interleaves the generation of CoT with knowledge retrieval from corresponding KGs, iteratively guiding both retrieval and reasoning for multi-step questions. MindMap (Wen et al., 2023) introduce a plug-and-play approach to evoke graph-of-thoughts reasoning in LLMs. Similarly, RoG (Luo et al., 2023b) use KGs to create faithful reasoning paths based on various relations, enabling interpretable reasoning in LLMs.
@@
Appendix D Prompt List
@@
D.1 Knowledge-Augmented Generation